# Dance-to-Music Generation with Encoder-based Textual Inversion of Diffusion Models

### v1

@@ 4.5 Ablation Study
We first investigate the role of genre encoder that engage into dance-to-music generation. We introduce a variant called CLAP that replaces genre labels with audio embeddings derived from other segments of the same audio [50]. When the genre label is replaced with CLAP audio embeddings, experimental results indicate a decrease in all metrics. This observation indicates that the highly repetitive or monotonous audios in the AIST++ dataset is not suitable for facilitating the genre encoder in mapping CLAP audio embeddings to the textual space. As a result, achieving example-based genre control becomes challenging.
@@
We design two projectors to map rhythm sequences to the textual space: “MLP”, which utilizes a combination of multiple linear layers and activation layers as the projector, and “Attn+Pos”, which incorporates position embedding and attention modules as the projector. The results, as shown in Table  2, demonstrate that the Riffusion-based experiments perform better with the former projector and poorly with the latter. Conversely, the MUSICGEN-based experiments exhibit better performance with the latter projector. This indicates that MUSICGEN is more sensitive to position embedding.
4.6 Discussions and Limitations
Our approach generates diverse and high-quality music that aligns with videos, allowing for personalized, dynamic, and interactive music experiences in the context of dance videos. Furthermore, our method can be extended to various other physical activities (e.g., “jump rope”, “martial arts”, and “artistic gymnastics”). For limitation, our method currently supports only fixed-length video segments for music composition. The flexibility of accommodating variable-length segments would enhance the applicability of our method in real-world scenarios.
5 Conclusion
In this paper, we introduce an encoder-based textual inversion technique to seamlessly integrate rhythm and genre control into pre-trained text-to-music models. This approach offers a plug-and-play solution for text-to-music models. We further introduce an enhanced rhythm alignment evaluation metric and conduct a comprehensive assessment encompassing rhythm, audio quality, and genre. Experimental results demonstrate that our method can generate high-quality music that is synchronized with the beats across various genres, even with a limited dataset of 50 songs. Furthermore, our approach seamlessly adapts to in-the-wild data and enables flexible editing of high-level music attributes using text.
Future research may delve into the integration of multimodal motion information to achieve more robust beat generation results. Additionally, incorporating example-based genre control could provide more flexible user experiences in dance-to-music generation.

### v2

@@ 4.5. Ablation Study
We first investigate the role of genre encoder that engage into dance-to-music generation. We introduce a variant called CLAP that replaces genre labels with audio embeddings derived from other segments of the same audio (Wu et al., 2023a). When the genre label is replaced with CLAP audio embeddings, experimental results indicate a decrease in all metrics except TD. This observation indicates that the highly repetitive or monotonous audios in the AIST++ dataset is not suitable for facilitating the genre encoder in mapping CLAP audio embeddings to the textual space. As a result, achieving example-based genre control becomes challenging.
Table 5. Ablation study of our method. “M” and “R” respectively represents “MUSICGEN” and “Riffusion”. “GL” represents genre label. “CLAP” represents utilization of CLAP audio embeddings to replace genre labels. “MLP” represents the adoption of a Multi-Layer Perceptron (MLP) as the projector in the rhythm encoder. “Attn+Pos” denotes the employment of attention mechanism and positional encoding as the projector in the rhythm encoder.
Base Encoder Rhythm Quality Genre
BCS $\uparrow$ BHS $\uparrow$ F1-score $\uparrow$ TD $\downarrow$ FAD $\downarrow$ CLAP $\uparrow$
M CLAP 0.3925 0.3155 0.3498 13.47 6.821 0.4588
MLP + GL 0.3958 0.3643 0.3794 15.51 6.469 0.4695
Attn + Pos 0.4118 0.3874 0.3992 16.06 6.014 0.4685
R CLAP 0.4480 0.3085 0.3654 14.49 4.611 0.6477
MLP + GL 0.4761 0.4398 0.4572 20.34 3.416  0.7680
Attn + Pos 0.3806 0.2768 0.3205 21.69 3.814 0.7524
@@
We design two projectors to map rhythm sequences into the textual space: “MLP”, which utilizes a combination of multiple linear layers and activation layers as the projector, and “Attn+Pos”, which incorporates position embedding and attention modules as the projector. As indicated in Table 5, experiments using the Riffusion framework demonstrate improved performance with the MLP projector but diminished results with the Attn+Pos projector. In contrast, the experiments conducted with MUSICGEN show enhanced performance with the Attn+Pos projector, suggesting that MUSICGEN is more sensitive to position embeddings.
4.6. Discussions and Limitations
Our approach generates diverse and high-quality music that aligns with videos, allowing for personalized and interactive music experiences in the context of dance videos. Furthermore, our method can be extended to in-the-wild data (e.g., “film”, “jump rope” and “artistic gymnastics”). For limitation, our method currently supports only fixed-length video segments for music composition. The flexibility of accommodating variable-length segments would enhance the applicability of our method in real-world scenarios.
5. Conclusion
This paper introduces an encoder-based textual inversion technique to seamlessly integrate rhythm and genre control into pre-trained text-to-music models. Our approach offers a plug-and-play solution for text-to-music models. We conduct a comprehensive evaluation on two datasets, encompassing rhythm, audio quality, and genre. Experimental results demonstrate that our method can generate high-quality music that is synchronized with the videos across various genres. Furthermore, our approach adapts to changes in tempo and enables flexible editing of high-level music attributes using text prompts. Future research may delve into the integration of multimodal motion information to achieve more robust beat generation results. Additionally, incorporating example-based genre control could provide more flexible user experiences in dance-to-music generation.
Acknowledgements.
We thank Xue Song for preparing some dance video data and Minyan Luo for selection of the results. This work was supported in part by the National Natural Science Foundation of China under nos. U20B2070 and 62102162, in part by the Beijing Science and Technology Plan Project under no. Z231100005923033, in part by the National Science and Technology Council under no. 111-2221-E-006-112-MY3, Taiwan, in part by the German Research Foundation (DFG) Project under no. 508324734, and in part by Kuaishou.